OpenStudy (anonymous):

f(x)=2x^2e^5x+1 find derivative

OpenStudy (anonymous):

@Mcjay Are you familiar with the chain rule? And welcome to OpenStudy :-D

OpenStudy (anonymous):

Chain Rule: for: \[u^n \] then: \[(n*u^{n-1})*\frac{d}{du}u\] Do the power rule, then take the derivative of what's inside the "shell"/"layer"/"whatever"

OpenStudy (anonymous):

basically i'm looking how i can getlocal max and min

OpenStudy (anonymous):

after derivative i find this i dont know if is correct 20 x e^5x

OpenStudy (anonymous):

Local max & min will occur at critical points. To find critical points you need to find the first derivative, then find what the input needs to be to make the output (the slope) be zero (be horizontal). f(x) = y f'(x) = y' = \(\large\frac{dy}{dy}\) When y' = 0 , you'll have the critical points. Write up what you think the derivative is into the equation editor (click the "\(\Sigma\) Equation" button below) , you can write fractions by doing this: frac{dy}{dx} turns into \(\large\frac{dy}{dy}\)

OpenStudy (anonymous):

\[f(x) = 2 x^2 e^5 x+1\] \[f'(x)=\frac{d}{dx}(2 x^2 e^5 x+1)\] \[f'(x)=2 e^5 (\frac{d}{dx}(x^3))+\frac{d}{dx}(1)\] \[f'(x)=2 e^5 (3x^2)+0 \ \ \ \ \ dx\] \[f'(x)=6 e^5 x^2 \ \ \ \ \ \ dx\]

OpenStudy (anonymous):

\[dy/dx 2x^2e ^(5x) +1

OpenStudy (anonymous):

5x

OpenStudy (anonymous):

Actually I just noticed I had a typo, the derivative of y is dy, so y' = dy, not dy/dy. Derp.

OpenStudy (anonymous):

right

OpenStudy (anonymous):

Minor detail, but just to point that out if you're writing this down for your notes.

OpenStudy (anonymous):

i try to write e ^(5x) but i dont get it

OpenStudy (anonymous):

you can divide both sides by dx, to just get dy/dx which of course IS your instantaneous slope. Average slope: \[m=\frac{\triangle y}{\triangle x}=\frac{y_2-y_1}{x_2-x_1}\] Instantaneous slope: \[m=\frac{dy}{dx}\]

OpenStudy (anonymous):

Oh you mean what you intended to write wasn't \(2x^2e^5x+1\)?

OpenStudy (anonymous):

was not e is exponent 5x

OpenStudy (anonymous):

Yeah you REALLY got to watch your parenthesis here, especially when asking question online. We can't see your original problem of course, so you got to type it in to make sure you communicate clearly. \[\frac{d}{dx}(2 x^2 e^{5 x}+1) \ ????\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

sorry about that it is my first time i use online help

OpenStudy (anonymous):

help please

OpenStudy (anonymous):

Factor out the two, apply the derivative to both terms (it's associative): \[2 (\frac{d}{dx}(e^{5 x} x^2))+\frac{d}{dx}(1)\] Using the product rule, "first d second + second d first": \[2(x^2 (\frac{d}{dx}(e^{5 x}))+e^{5 x} (\frac{d}{dx}(x^2)))\] Using the chain rule: \[2(x^2 (e^{5 x} (\frac{d}{dx}(5 x)))+e^{5 x} (\frac{d}{dx}(x^2)))\]

OpenStudy (anonymous):

Can you finish from here? You'll have this mess simplify to something like this (if you factor out the common term): \[2xe^{5 x}(5 x+2) = \frac{dy}{dx} = m = slope\] You need to find what makes the slope zero. What the "x" input needs to be.

OpenStudy (anonymous):

5x+2 =0

OpenStudy (anonymous):

Yep, that's one of the roots, now what's the other? :-)

OpenStudy (anonymous):

x=-2/5

OpenStudy (anonymous):

Anything times zero is...?

OpenStudy (anonymous):

is 0

OpenStudy (anonymous):

i think to find local min and max u find x first

OpenStudy (anonymous):

Yep, that's the other. Kind of hard to see in the original graph but that's the two times it reaches zero.